# Geometry and Perception Guided Gaussians for Multiview-consistent 3D Generation from a Single Image

### v2

@@ Abstract
Index Terms:
Single-view 3D Generation, Geometry, Perception, Prior Knowledge, 3D Gaussian Splatting, Consistency
@@ I Introduction
Single-view 3D generation is a basic task in computer vision, which aims to predict the full 3D geometry (e.g., a mesh, point cloud) of an object or a scene from an input 2D image [1, 2, 3, 4, 5, 6]. It requires that the resulting 3D content not only appears natural and consistent from all viewpoints, but also accounts for multiple plausible interpretations of unseen regions. Humans possess a remarkable ability to envision a complete 3D structure from a single image, even when certain parts of the object remain hidden. This gives credits to prior knowledge of familiar objects, so that the human brain infers and represents missing regions to form a coherent 3D concept. In the fields of computer vision and graphics, substantial progress has been made in deriving implicit 3D representations from images [7, 8, 9]. However, explicitly reconstructing 3D objects directly from 2D images remains a formidable challenge.
@@ II-C Single-view 3D Generation
Single-view 3D generation must yield fully detailed and interactive 3D content rather than only multiview images. A complete 3D model is essential for applications in virtual reality, augmented reality and 3D printing, all of which require accurate depth, texture and geometry. Accordingly, the goal is to produce high-fidelity 3D models that capture fine-grained details, preserve structural consistency and support seamless manipulation and rendering from any viewpoint. Achieving this objective demands both improved consistency across views and innovations in 3D reconstruction techniques, for example implicit neural representations such as NeRF or explicit mesh-based methods. We argue that direct point cloud/voxel diffusion models and Gaussian splatting approaches are inherently superior for single-view 3D generation.
@@ III-B Different Prior Knowledge Guidance
Based on the preliminaries, we analyze the role of different priors in improving generation quality. Specifically, we address challenges related to initializing and optimizing Gaussian representations. Unlike reconstruction tasks, structure-from-motion (SfM) methods cannot be applied due to the only one input image. Additionally, optimization becomes challenging given only a frontal viewpoint, limiting the guidance available for Gaussian training.
@@ III-B1 Geometry Prior
We assume that geometry prior provides the initialization information for the single input image. Previous work [42] has shown that extending the geometric representation can achieve alignment between the geometrical data and the CLIP embedding space. The geometry prior can be acquired through retrieving structurally analogous 3D shapes from an external dataset, following [43], which chooses to incorporate external 3D prior information into text-to-3D generation by designing a shape supervision loss based on prior 3D shapes. However, it is crucial to recognize the inherent distinction between these two generation tasks: while text-to-3D generation emphasizes high-level semantic alignment with text descriptions, image-to-3D generation requires that the frontal perspective of the generated object is exactly the same as the provided image in pixel level. Directly applying 3D shape supervision loss may lead to conflicts between the input perspective and the generated perspective, so we adopt the retrieved 3D shapes to initialize Gaussian positions so as to ensure the coarse shape representation, which makes it easier to converge to the target shape after optimization.
@@ III-B2 Perception Prior
Following [45], we consider to address this limitation by incorporating depth information. Specifically, we utilize Marigold [45] to estimate rough depth maps from both the input and generated multiview images. We combine them to form an coarse point cloud to initialize the 3D Gaussian Splatting model, which provides a good solution for initialization issues.
@@ III-C1 Perceptual Augmentation Loss
Specifically, we denote that the input image is $\tilde{\mathbf{I}}_{R}^{0}$ with a foreground mask $\tilde{\mathbf{I}}_{A}^{0}$, the generated $i$-th multiview image and its foreground mask are $\tilde{\mathbf{I}}_{R}^{i}$ and $\tilde{\mathbf{I}}_{A}^{i}$, (we utilize the Rembg library [46] for isolating foreground objects to obtain $\tilde{\mathbf{I}}_{A}^{i}$.) where $i>0$ . for each 3D Gaussian branch, we optimize the reference view image $\mathbf{I}_{R}^{0}$ and mask $\mathbf{I}_{A}^{0}$ to align with the input:
@@
In addition to the mean-square-error, we add a D-SSIM [47] term to promote the structural similarity:
@@ III-C2 Stable SDS Loss
Eq. 8 admits an interpretation through the lens of contrastive learning. Specifically, positive samples are critical in contrastive learning as they often share similar latent space. In our task, images rendered from different Gaussian branches under the same camera pose $\mathbf{p}$ are regarded as positive samples, which have the same conditions (the input image and the same camera pose), and after $l$ iterations, the images rendered from the different Gaussian branches are similar at the pixel level. We then analyze the three components of different Gaussian branches $i$ and $j$. We utilize alphabetic superscripts to differentiate among Gaussian branches. For a camera pose $\mathbf{p}$, we have $\delta_{\text{cond}}^{i}\approx\delta_{\text{cond}}^{j}$ and $\delta_{\text{real}}^{i}\approx\delta_{\text{real}}^{j}$ after $l$ iterations. This derivation holds under the condition that both components reach convergence within $l$ iterative steps. For bias $\delta_{\text{bias}}$, $\delta_{\text{bias}}^{i}\neq\delta_{\text{bias}}^{j}$ because of its stochasticity. The term $)(\epsilon_{\phi}(\mathbf{z}_{t}^{i},...)-\epsilon_{\phi}(\mathbf{z}_{t}^{j},$...)) in Eq. 8 can be approximated as $(\delta_{\text{bias}}^{i}-\delta_{\text{bias}}^{j})$. The term $(\epsilon_{\phi}(\mathbf{z}_{t}^{i},...)-\epsilon)$ can be approximated as $\delta_{\text{bias}}$. So we rewrite Eq. 8 as:
@@ III-C3 Reprojection Loss for Depth Consistency
where $\text{D}_{2}$ is the depth image rendered by 3D Gaussians in view 2 and $|\text{D}_{2}|$ means the total pixels in the depth image. The process of calculating the reprojection loss is shown in Fig. 4.
@@ IV Experiments
We test the proposed method on the Google Scanned Object dataset (in Sec. IV-A) for novel view synthesis task (in Sec. IV-B) and single view reconstruction (in Sec. IV-C) task without training any neural networks. We show that our results achieve more accurate view synthesis as well as yield better 3D reconstruction compared to previous lightweight approaches. We also present ablation studies (in Sec. IV-D) on crucial modules mentioned in Sec. III-C.
@@ IV-A2 Implementation Details
We train 1200 steps for each object and set $l=600$. Sec. III-D clarified the initialization of first and second Gaussian branches. For the third branch, we sample points inside a sphere of radius 0.5. The number of points is 10000 for the three initializations. For the first branch, we introduce mesh sampling and farthest point sampling (FPS) to downsample the original 3D object to 10000 points. For the second branch, we introduce FPS to downsample to 10000 points.
@@ IV-A3 Baseline
For novel view synthesis task, we adopt Zero123 [37], RealFusion [6], SyncDreamer [18], DreamGaussian [1] and MVD-Fusion [2] as baseline methods. Given an input image of an object and target view parameters, Zero123 can estimate the image of the target view. RealFusion is based on Stable Diffusion [16] and the SDS loss for single-view reconstruction. SyncDreamer and MVDFusion can generate multiview images simultaneously based on a fine-turned 2D diffusion model. DreamGaussian adopts 3D Gaussian Splatting as its 3D representation and only uses a input image and the SDS loss to optimize the object without training additional models.
For single view reconstruction task, we adopt Zero123 [37], RealFusion [6], Magic123 [5], One-2-3-45 [3], Point-E [15], Shap-E [11], DreamGaussian [1], SyncDreamer [18] and MVDFusion [2] as baseline methods. Magic123 combines Zero123 with RealFusion to further improve the quality of the reconstruction. One-2-3-45 directly regresses the sign distance function from the output images of Zero123. Point-E and Shap-E are 3D generative models trained on a large internal OpenAI 3D dataset, both of which are able to convert a single-view image into a point cloud or a shape encoded in an MLP. For Point-E, we convert the generated point clouds to SDFs for shape reconstruction. We should note that the diffusion-based methods require additional methods to render or distill objectives. e.g. Zero-1-to-3 requires an additional SDS distillation to generate 3D shapes whereas SyncDreamer relies on training Neus or NeRF.
@@ IV-A4 Metrics
For novel view synthesis, we adopt commonly used metrics: PSNR, SSIM [47], and LPIPS [48]. For single view reconstruction, We report the commonly used Chamfer Distances (CD) and Volume IoU between ground-truth shapes and reconstructed shapes.
@@ IV-B Novel View Synthesis
For this task, the quantitative results are shown in Table II and the qualitative results are shown in Fig. 5. RealFusion [6] applies a NeRF model to distill the stable diffusion model, but is unable to produce visually plausible images. Zero123 [37] produces visually reasonable images but suffers from inconsistency and instability among the generated images. SyncDreamer [18] and MVD-Fusion [2] are limited to specific perspective. Meanwhiles, MVD-Fusion is prone to perspective inaccuracy and multiple perspectives inconsistency, as shown in the third column of Fig. 5. DreamGaussian [1] faces oversaturated textures or blurry back-view textures as shown in the second and fifth rows of Fig. 5. Our method fully utilizes geometry and perception prior knowledge to supervise different Gaussian branches, achieving competitive generation quality.
@@ IV-C Single-view Reconstruction
For this task, we show the quantitative results in Table III and the qualitative comparison in Fig 6. Magic123 [30] heavily relies on the estimated depth values of the input view, leading to incorrect results and instability. Point-E [15] and Shap-E [11] tend to produce incomplete meshes. One-2-3-45[3] reconstructs meshes from the outputs of Zero123, but the reconstruction suffers from loss of details due to image quality constraints. SyncDreamer [18] uses vanilla NeuS [9] for reconstruction. The method is incapable of direct 3D mesh generation. Even with other methods, its texture quality still cannot reach the ideal standard. MVD-Fusion [2] can directly obtain the point cloud but cannot generate meshes. DreamGaussian [1] generate imcomplete 3D geometric structure in some cases, as shown in the third row of Fig. 6. Wonder3D [4] produces preliminary geometric approximations of objects, yet exhibits limitations in fine-grained detail reproduction and occasional inaccuracies in 3D structural reconstruction. Comparative evaluations demonstrate that our approach significantly improves reconstruction fidelity while simultaneously maintaining surface continuity while preserving high-frequency geometric features.
Furthermore, we compare the consumption of different methods in single-view reconstruction task, as shown in Table. IV. MVD-Fusion, Zero123, SyncDreamer and Wonder3D are based on fine-turning diffusion models to improve generation consistency, which leads to a considerable amount of time and GPUs consumption. Meanwhile, althougn MVD-Fusion exhibits rapid inference, it can only generate point clouds instead of meshes without other tools. Although DreamGaussian achieves notable improvements in computational efficiency and resource utilization over prior approaches, limitations persist in the quality of its output textures and geometric structures, suggesting avenues for future enhancement. Our method achieve high-quality generation without the need to train additional neural networks, while also maintaining low computational cost during inference.
@@ IV-D1 Different Branches
To show how different Gaussian branches affect the result and the improvement of interactions in branches, we conduct experiments with only one of the three Gaussian branches without interactions (e.g. the stable SDS). As we can see in Table V, only one branch shows a decrease in three indicators, demonstrating the necessity of the interaction of different branches.
@@ IV-D2 Different Modules
To show how different modules mentioned in Sec. III-C affect the generation. We discard perceptual augmentation loss, stable SDS loss, and reprojection loss respectively. As we can see in Table VI, the absence of any module will lead to a deterioration of quantitative results. We also display the novel view synthesis results based on the absence of the proposed modules. As we can see in Fig. 7, the absence of reprojection loss lead to local noise, while the absence of stable SDS loss and perception augmentation loss lead to blurry or oversaturated textures. The results with full modules exhibit reduced blurriness and improved accuracy.
@@ V Conclusion
We propose a novel method for reconstructing detailed 3D objects from a single image. Our method employs Gaussian splatting, guided by geometry and perception priors, to address the challenges of high memory requirements. To ensure effective knowledge transfer, we introduce a stable SDS loss for fine-grained prior distillation. Furthermore, we propose a reprojection-based strategy to reinforce geometric consistency in depth estimation. Experimental results demonstrate that our approach achieves superior generation fidelity, effectively improving the multiview consistency and geometric details loss.

### v3

@@ I Introduction
Single-view 3D generation is a basic multi-modality task in multimedia, which aims to predict the full 3D geometry (e.g., a mesh, point cloud) of an object or a scene from an input 2D image [1, 2, 3, 4, 5, 6, 7, 8, 9]. It requires that the resulting 3D content not only appears natural and consistent from all viewpoints, but also accounts for multiple plausible interpretations of unseen regions. Humans possess a remarkable ability to envision a complete 3D structure from a single image, even when certain parts of the object remain hidden. This gives credits to prior knowledge of familiar objects, so that the human brain infers and represents missing regions to form a coherent 3D concept. In the fields of computer vision and graphics, substantial progress has been made in deriving implicit 3D representations from images [10, 11, 12]. However, explicitly reconstructing 3D objects directly from 2D images remains a formidable challenge.
@@ II-C Single-view 3D Generation
Single-view 3D generation must yield fully detailed and interactive 3D content rather than only multiview images. A complete 3D model is essential for applications in virtual reality, augmented reality and 3D printing, all of which require accurate depth, texture and geometry. Accordingly, the goal is to produce high-fidelity 3D models that capture fine-grained details, preserve structural consistency and support seamless manipulation and rendering from any viewpoint. Achieving this objective demands both improved consistency across views and innovations in 3D reconstruction techniques. We argue that direct point cloud/voxel diffusion models and Gaussian splatting approaches are inherently superior for single-view 3D generation.
@@ III-B Different Prior Knowledge Guidance
We introduce the proposed method by first analyzing the role of different priors in improving generation quality. Specifically, we address challenges related to initializing and optimizing Gaussian representations. Unlike reconstruction tasks, structure-from-motion (SfM) methods cannot be applied due to the only one input image. Additionally, optimization becomes challenging given only a frontal viewpoint, limiting the guidance available for Gaussian training.
@@ III-B1 Geometry Prior
We assume that geometry prior provides the initialization information for the single input image. A previous work [45] has shown that extending the geometric representation can achieve alignment between the geometrical data and the CLIP embedding space. The geometry prior can be acquired through retrieving structurally analogous 3D shapes from an external dataset, following [46], which chooses to incorporate external 3D prior information into text-to-3D generation by designing a shape supervision loss based on prior 3D shapes. However, it is crucial to recognize the inherent distinction between these two generation tasks: while text-to-3D generation emphasizes high-level semantic alignment with text descriptions, image-to-3D generation requires that the frontal perspective of the generated object is exactly the same as the provided image in pixel level. Directly applying 3D shape supervision loss may lead to conflicts between the input perspective and the generated perspective, so we adopt the retrieved 3D shapes to initialize Gaussian positions so as to ensure the coarse shape representation, which makes it easier to converge to the target shape after optimization.
@@ III-B2 Perception Prior
Following [48], we consider to address this limitation by incorporating depth information. Specifically, we utilize Marigold to estimate rough depth maps from both the input and generated multiview images. We combine them to form a coarse point cloud to initialize the 3D Gaussian Splatting model, which provides a good solution for initialization issues.
@@ III-C1 Perceptual Augmentation Loss
Specifically, we denote that the input image is $\tilde{\mathbf{I}}_{R}^{0}$ with a foreground mask $\tilde{\mathbf{I}}_{A}^{0}$, the generated $i$-th multiview image and its foreground mask [^1] We utilize the Rembg library [49] for isolating foreground objects to obtain the foreground mask. are $\tilde{\mathbf{I}}_{R}^{i}$ and $\tilde{\mathbf{I}}_{A}^{i}$, where $i>0$. For each 3D Gaussian branch, we optimize the reference view image $\mathbf{I}_{R}^{0}$ and mask $\mathbf{I}_{A}^{0}$ to align with the input:
@@
In addition to the mean-squared-error, we add a D-SSIM [50] term to promote the structural similarity:
@@ III-C2 Stable SDS Loss
Eq. 8 admits an interpretation through the lens of contrastive learning. Specifically, positive samples are critical in contrastive learning as they often share similar latent space. In our task, images rendered from different Gaussian branches under the same camera pose $\mathbf{p}$ are regarded as positive samples, which have the same conditions (the input image and the same camera pose), and after $l$ iterations, the images rendered from the different Gaussian branches are similar at the pixel level. We then analyze the three components of different Gaussian branches $i$ and $j$. We utilize alphabetic superscripts to differentiate among Gaussian branches. For a camera pose $\mathbf{p}$, we have $\delta_{\text{cond}}^{i}\approx\delta_{\text{cond}}^{j}$ and $\delta_{\text{real}}^{i}\approx\delta_{\text{real}}^{j}$ after $l$ iterations. This derivation holds under the condition that both components reach convergence within $l$ iterative steps. For bias $\delta_{\text{bias}}$, $\delta_{\text{bias}}^{i}\neq\delta_{\text{bias}}^{j}$ because of its stochasticity. The term $)(\epsilon_{\phi}(\mathbf{z}_{t}^{i},...)-\epsilon_{\phi}(\mathbf{z}_{t}^{j},$...)) in Eq. 8 can be approximated as $(\delta_{\text{bias}}^{i}-\delta_{\text{bias}}^{j})$. The term $(\epsilon_{\phi}(\mathbf{z}_{t}^{i},...)-\epsilon)$ can be approximated as $\delta_{\text{bias}}$. So we approximate Eq. 8 as:
@@ III-C3 Reprojection Loss for Depth Consistency
where $\text{D}_{1}$ and $\text{D}_{2}$ are the depth images rendered by 3D Gaussians in view 1 and view 2, respectively, and $|\text{D}_{2}|$ means the total pixels in the depth image. The process of calculating the reprojection loss is shown in Fig. 4.
@@ IV Experiments
We test the proposed method on the Google Scanned Object dataset (introduced in Sec. IV-A) for the novel view synthesis task (in Sec. IV-B) and single view reconstruction (in Sec. IV-C) without training any neural network. We show that our results achieve more accurate view synthesis as well as yield better 3D reconstruction compared to previous lightweight approaches. We also present ablation studies (in Sec. IV-D) on crucial modules mentioned in Sec. III-C.
@@ IV-A2 Implementation Details
We train 1200 steps for each object and set $l=600$. Sec. III-D clarified the initialization of first and second Gaussian branches. For the third branch, we sample points inside a sphere of radius 0.5. The number of points is 10k for the three initializations. For the first branch, we introduce mesh sampling and farthest point sampling (FPS) to downsample the 3D object to 10k points. For the second branch, we introduce FPS to downsample to 10k points.
@@ IV-A3 Baseline
For the novel view synthesis task, we adopt Zero123 [40], RealFusion [6], SyncDreamer [21], DreamGaussian [1] and MVD-Fusion [2] as baseline methods. Given an input image of an object and target view parameters, Zero123 can estimate the image of the target view. RealFusion is based on Stable Diffusion [19] and the SDS loss for single-view reconstruction. SyncDreamer and MVDFusion can generate multiview images simultaneously based on a fine-turned 2D diffusion model. DreamGaussian adopts 3D Gaussian Splatting as 3D representation and only uses an input image and the SDS to optimize the object without training additional models.
For single view reconstruction, we adopt Zero123 [40], RealFusion [6], Magic123 [5], One-2-3-45 [3], Point-E [18], Shap-E [14], DreamGaussian [1], SyncDreamer [21] and MVDFusion [2] as baseline methods. Magic123 combines Zero123 with RealFusion to further improve the quality of the reconstruction. One-2-3-45 directly regresses the sign distance function from the output images of Zero123. Point-E and Shap-E are 3D generative models trained on a large internal OpenAI 3D dataset, both of which are able to convert a single-view image into a point cloud or a shape encoded in an MLP. For Point-E, we convert the generated point clouds to SDFs for shape reconstruction. We should note that the diffusion-based methods require additional methods to render or distill objectives. For instance, Zero-1-to-3 requires an additional SDS distillation to generate 3D shapes whereas SyncDreamer relies on training Neus or NeRF.
@@ IV-A4 Metrics
For novel view synthesis, we adopt commonly used metrics: PSNR, SSIM [50], and LPIPS [51]. For single view reconstruction, we report the commonly used Chamfer Distance (CD) and Volume IoU between ground-truth shapes and reconstructed shapes.
@@ IV-B Novel View Synthesis
For this task, the quantitative results are shown in Table II and the qualitative results are shown in Fig. 5. RealFusion [6] applies a NeRF model to distill the stable diffusion model, but is unable to produce visually plausible images. Zero123 [40] produces visually reasonable images but suffers from inconsistency and instability among the generated images. SyncDreamer [21] and MVD-Fusion [2] are limited to specific perspectives. Meanwhiles, MVD-Fusion is prone to perspective inaccuracy and multiple-perspective inconsistency, as shown in the third column of Fig. 5. DreamGaussian [1] faces oversaturated textures or blurry back-view textures as shown in the second and fifth rows of Fig. 5. Our method utilizes geometry and perception prior knowledge to supervise different Gaussian branches, achieving competitive generation quality.
@@ IV-C Single-view Reconstruction
For this task, we show the quantitative results in Table III and the qualitative comparison in Fig 6. Magic123 [33] heavily relies on the estimated depth values of the input view, leading to incorrect results and instability. Point-E [18] and Shap-E [14] tend to produce incomplete meshes. One-2-3-45 [3] reconstructs meshes from the outputs of Zero123, but the reconstruction suffers from loss of details due to image quality constraints. SyncDreamer [21] uses vanilla NeuS [12] for reconstruction. The method is incapable of direct 3D mesh generation. Even with other methods, its texture quality still cannot reach the ideal standard. MVD-Fusion [2] can directly obtain the point cloud but cannot generate meshes. DreamGaussian [1] generates imcomplete 3D geometric structure in some cases, as shown in the third row of Fig. 6. Wonder3D [4] produces preliminary geometric approximations of objects, yet exhibits limitations in fine-grained detail reproduction and occasional inaccuracies in 3D structural reconstruction. Comparative evaluations demonstrate that our approach significantly improves reconstruction fidelity while simultaneously maintaining surface continuity and preserving high-frequency geometric features.
Furthermore, we compare the consumption of different methods in the single-view reconstruction task, as shown in Table IV. MVD-Fusion, Zero123, SyncDreamer and Wonder3D are based on fine-turning diffusion models to improve generation consistency, which leads to a considerable amount of time and GPU consumption. Meanwhile, although MVD-Fusion exhibits rapid inference, it can only generate point clouds instead of meshes without other tools. Although DreamGaussian achieves notable improvements in computational efficiency and resource utilization over prior approaches, limitations persist in the quality of its output textures and geometric structures, suggesting avenues for future enhancement. Our method achieves high-quality generation without the need to train additional neural networks, while also maintaining low computational cost during inference.
@@ IV-D1 Different Branches
To show how different Gaussian branches affect the result and the improvement of interactions in branches, we conduct experiments with only one of the three Gaussian branches without interactions (e.g., the stable SDS). As we see in Table V, only one branch shows decrease in three indicators, demonstrating the necessity of the interaction of different branches.
@@ IV-D2 Different Modules
To show how different modules mentioned in Sec. III-C contribute to the generation results, we discard perceptual augmentation loss, stable SDS loss, and reprojection loss, respectively. As we see in Table VI, the absence of any module will lead to a deterioration of quantitative results. We also display the novel view synthesis results based on the absence of the proposed modules. As in Fig. 7, the absence of reprojection loss leads to local noise, while the absence of stable SDS loss and perception augmentation loss leads to blurry or oversaturated textures. The results with full modules exhibit reduced blurriness and improved accuracy.
@@ V Conclusion
We propose a novel method for reconstructing detailed 3D objects from a single image. Our method employs Gaussian splatting, guided by geometry and perception priors, to address the challenges of high memory requirements. To ensure effective knowledge transfer, we introduce a stable SDS loss for fine-grained prior distillation. Furthermore, we propose a reprojection-based strategy to reinforce geometric consistency in depth estimation. Experimental results demonstrate that our approach achieves superior generation fidelity, effectively improving the multiview consistency and geometric detail.